OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y=x^2-1 ; y=-x-1 . PLEASE HELP !

OpenStudy (anonymous):

-x-1=x^2-1 x^2+x=0 x(x-1)=0 x=0 x=1